OpenStudy (kirbykirby):

Misprints are distributed randomly and uniformly in a book at a rate of 2 per 100 lines. What is the probability that a line is free of misprints?

OpenStudy (kirbykirby):

I thought this was simply saying that 2 lines of 100 have errors So, there are 100-2=98 lines out of 100 that are error free (on avg. ) So, Probability = 98/100 = 0.98 But, my book uses a Poisson distribution: Let X be the # of errors μ=λt where λ = 0.02/line, t = 1 line, so μ=0.02 \[P(X=0)=\frac{(0.02)^0e^{−0.02}}{0!}\] =0.9802 (which is still close to 0.98... but I wonder why we need the Poisson here?)

OpenStudy (amistre64):

randomly and uniformily, so its not a normal distribution

OpenStudy (amistre64):

Poison does seem most applicable

OpenStudy (amistre64):

the simplicity might be so that you can reasonably determine the correctness of the poison method for this

OpenStudy (kirbykirby):

Hm. So you can use the Poisson for things that don't involve periods of time (which is what we usually saw)?

OpenStudy (amistre64):

correct. most likely this problem can relate very well to a time based setup

OpenStudy (kirbykirby):

oh ok interesting. thanks :)